Calculate the partial pressures of all components in a mixture that contain 1.0 mole of CO2, 2.0 moles CH4 and 1.5 moles of CO in a constant volume vessel at 120.0 kPa and 20oC.




1
Expert's answer
2021-10-27T06:41:20-0400

Compute for the total number of moles of gas.


nT = n CO2 + n CH4 + n CO


nT = 1.0 mol + 2.0 mol + 1.5 mol


nT = 4.5 mol


Compute the mole fraction of each gas.


For CO2


X CO2 = n CO2 / nT


X CO2 = 1.0 mol / 4.5 mol


X CO2 = 0.22222


For CH4


X CH4 = n CH4 / nT


X CH4 = 2.0 mol / 4.5 mol


X CH4 = 0.44444


For CO


X CO = n CO / nTX CO = 0.33333


Calculate the partial pressures of each ga


For C


P CO2 = (X CO2)(P


P CO2 = (0.22222)(182.34 kP


P CO2 = 40.52 k


For C


P CH4 = (X CH4)(P


P CH4 = (0.44444)(182.34 kP


P CH4 = 81.04 k


For


P CO = (X CO)(P


P CO = (0.33333)(182.34 kP


P CO = 60.78 kPaa)T)COPaa)T)H4Paa)T)O2s.